AnimeGhoul8863:

AM I RIGHT? A carton of milk has spilled on a tile floor. The milk flow can be expressed with the function m(t) = 10t, where t represents time in minutes and m represents how far the milk is spreading. The flowing milk is creating a circular pattern on the tile. The area of the pattern can be expressed as A(m) = πm2. Part A: Find the area of the circle of spilled milk as a function of time, or A[m(t)]. Show your work. (6 points) Part B: How large is the area of spilled milk after 3 minutes? You may use 3.14 to approximate π in this problem. (4 points)

AnimeGhoul8863:

Part A: \pi m^2 \\ \\ A = \pi (10t)^2 \\ \\ A = 100\pi t^2 Part B: A(3) = 100\pi (3^2) = 900 \pi

NOpanda14:

as far as I know yes

AnimeGhoul8863:

is there a way u can help me make the answer more answer friendly like i dont get the lines and that stuff

dude:

Latex: Part A: \(\pi m^2 \\ \\ A = \pi (10t)^2 \\ \\ A = 100\pi t^2\) Part B: \(A(3) = 100\pi (3^2) = 900 \pi\)

AnimeGhoul8863:

thx dude that helps alot